60kg man collide with 40kg man  with60N  what is the force experienced by the 40kg man

Grade:12th pass

1 Answers

Jaskirat singh
6 years ago
First we have to get the net accleration...So M=60kg and m=40kg, F=60NF=mass *accleration60=(60+40)*a60=100*a60/100=aa=0.6m/sNow for m=40kgF=m*aF=40*0.6F=24N